OpenStudy (anonymous):

the number n of aluminum cans used each year varies directly as the number of people p using cans. if 53 people use 18,815 cans in one yr how many cans are used in a city with 1,800.000

OpenStudy (anonymous):

53/18815=1800000/x 53x=18815*1800000 53x=33867000000 x=639 million cans